if Cos A+Cos 2 A=1 THEN FIND Sin 2 A+Sin 4 A=?

if Cos A+Cos2A=1 THEN FIND Sin2A+Sin4A=?

Grade:

4 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student, please find the solution:

Cos A+Cos2A=1
=> cosA = 1 – cos2A = sin2A

So, Required is
sin2A + sin4A
= Cos A+Cos2A
=1

I hope the solution is clear
Thanks.
Eshika gupta
5 years ago
Since cos A + cosA=1 
Now,cosA= 1-cos2A=sin2A
Then  taking L.H.S. sin2A+sin4A
=Sin2A(1+sin2A)
=CosA(1+cosA)
CosA+cos2A=1(given)
L.H.S =1 Hence proved
Rishab
5 years ago
Cosa+cos^2=1
Cosa+1-sin2a=1
Cosa-sin2a=0
Cosa=sin2a
Squaring both side 
Cos2a=sin4a
Sin2a+sin4a=1
Sin2a+cos2a=1
1=1
Rahul ram
4 years ago
CosA+cos^2=1
CosA+1-Sin^2A=1
CosA-Sin^2A=0
CosA=Sin^2A
Sq both side 
Cos^2A=Sin^4A
1-Sin^2A=Sin^4A
1=Sin^4A+Sin^2A
Hence ,
Ans is 1....